OpenStudy (anonymous):

can someone please help me with simplifying a radical with a nth root please

OpenStudy (mertsj):

Not unless we see the problem.

OpenStudy (anonymous):

\[\sqrt[4]{405}x ^{3}y ^{2}\]

OpenStudy (anonymous):

all under the fourth root sign

OpenStudy (mertsj):

Try to find a perfect fourth power that is a factor of 405

OpenStudy (mertsj):

They are 2^4=16 3^4=81. Try 81

OpenStudy (mertsj):